# Spatio-Temporal Graph Neural Network for Urban Spaces: Interpolating Citywide Traffic Volume

### v1

@@ Acknowledgements
For this set-up, we also provide the results of the ablation study (3). Since the setup in the main text uses a different train/validation/test split, direct comparisons are not meaningful.
@@
Appendix Appendix D Model Pipeline
@@
Appendix Appendix E Zero Inflated Negative Binomial and Negative Binomial Loss
@@
Appendix Appendix G True Zero Rate
@@
[…]
@@
Appendix Appendix H Strava Descriptive Statistics
@@
Appendix Appendix I Node Features
@@
Latitude and longitude street-level constant 2 numerical Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin (2023); OpenStreetMap contributors (2017) (Boeing, 2024)
@@
[…]
@@
[…]
@@
Degree see appendix Appendix J street-level constant 1 numerical computed based on Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin (2024a) computed based on (Boeing, 2024)
Betweenness see appendix Appendix J street-level constant 1 numerical computed based on Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin (2024a) computed based on (Boeing, 2024)
Closeness see appendix Appendix J street-level constant 1 numerical computed based on Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin (2024a) computed based on (Boeing, 2024)
Clustering coefficient see appendix Appendix J street-level constant 1 numerical computed based on Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin (2024a) computed based on (Boeing, 2024)
@@
$\sum$ 98 feature $\sum$ 79 features
Appendix Appendix J Connectivity Measures
@@
Appendix Appendix K Combination of Adjacency Matrices
@@
We only explain the adapted architecture of GNNUi (with a usage of ZINB loss), as only this case is employed here. However, the procedure for the other losses is the same and can be transferred accordingly. For both adjacency matrices, we seperately proceed with the same procedure as if we were to use only one adjacency matrix, passing the input through the three DGCN layers, using a relu activation function after the first two and a linear one after the last one to obtain $H_{3}\in\mathbb{R}^{n_{o}+n_{m}\times z}$. We then concatenate those along the second dimension to $H_{3},c\in\mathbb{R}^{n_{o}+n_{m}\times 2z}$. This is then passed through a ZINB layer to obtain $n$, $p$, and $\pi$. This is depicted in Figure 7.
@@
Table 6 presents the performance results of GNNUI, given the different combinations of adjacency matrices. Models that combine binary adjacency with another type perform well, likely benefiting from the strong performance of the binary structure. However, none of these combinations outperforms the standalone binary adjacency matrix, except for the KL for the taxi data (Binary is at 0.731 and Binary & Similarity at 0.744). Due to the small numerical difference, we think this is negligible. Interestingly, the combination of similarity and distance-based adjacency matrices improves upon the individual specifications in certain cases. For example, in New York, the Similarity + Distance Time combination achieves an MAE of 30.39, improving upon similarity (40.50) and distance time (31.68). While this suggests that hybrid approaches may be beneficial, they still fall short of the binary adjacency matrix.
@@
Appendix Appendix L Varying Percentage of Street Segments: For Similarity and Distance Based Adjacency Matrices

### v2

@@ Appendix C Temporal Generalization Ablation Study
For this set-up, we also provide the results of the ablation study (3). Since the setup in the main text uses a different train/validation/test split, direct comparisons between the observed errors are not meaningful.
@@
Appendix D Model Pipeline
@@
Appendix E Zero Inflated Negative Binomial and Negative Binomial Loss
@@ Appendix F Errors
Appendix G True Zero Rate
@@
$\text{true-zero rate}=\frac{\sum(\mathbbm{1}(T_{\text{t}}=0)\circ\mathbbm{1}(\mid\hat{T}_{\text{t}}\mid<\tau)\circ(1-M_{t})\circ N_{t})}{\sum(\mathbbm{1}(T_{\text{t}}=0)\circ(1-M_{t})\circ N_{t})}$ (23)
@@
Appendix H Strava Descriptive Statistics
@@
Appendix I Node Features
@@
Latitude and longitude street-level constant 2 numerical Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin (2023); OpenStreetMap contributors (2017) (Boeing, 2025)
@@
Type of bicycle lane $\in$ {no bicycle lane, bicycle lane on sidewalk, bicycle lane on street} street-level constant 1 categorical OpenStreetMap contributors (2017)
Type of street $\in$ {primary, residential, secondary, pedestrian, path, tertiary, unknown} street-level constant 1 categorical OpenStreetMap contributors (2017)
Type of street $\in$ {secondary, residential, primary, tertiary, motorway link, trunk, motorway, trunk link, tertiary link, primary link, secondary link, living street, unknown} street-level constant 1 categorical OpenStreetMap contributors (2017)
Street surface $\in$ { unknown, asphalt, sett, concrete, paving-stones, paved} street-level constant 1 categorical OpenStreetMap contributors (2017)
Number of car lanes $\in$ {unknown, 1, 2, $\cdots$} street-level constant 1 categorical OpenStreetMap contributors (2017)
@@
Max. permitted speed for cars in km/h, $\in$ {0, 30, 50} street-level constant 1 categorical OpenStreetMap contributors (2017)
Max. permitted speed for cars in mph, $\in$ {unknown, 20 mph or less, 15 mph, 20 mph, 25 mph, 30 mph, 35 mph, 40 mph, 45 mph, 50 mph} street-level constant 1 categorical OpenStreetMap contributors (2017)
Parking $\in$ {yes, no, unknown} street-level constant 1 categorical OpenStreetMap contributors (2017)
@@
Part of long distance cycling route street-level constant 1 numerical OpenStreetMap contributors (2017)
@@
Degree see appendix K street-level constant 1 numerical computed based on Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin (2024a) computed based on (Boeing, 2025)
Betweenness see appendix K street-level constant 1 numerical computed based on Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin (2024a) computed based on (Boeing, 2025)
Closeness see appendix K street-level constant 1 numerical computed based on Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin (2024a) computed based on (Boeing, 2025)
Clustering coefficient see appendix K street-level constant 1 numerical computed based on Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin (2024a) computed based on (Boeing, 2025)
@@
$\sum$ 98 features $\sum$ 79 features
Appendix J Feature Importance Analysis
Table 6: Feature Importance Groups for Taxi and Strava Datasets (values scaled by 10,000). Scores of one-hot encoded categorical features and features computed at multiple radii (e.g., hospital counts within 0.05–0.5 km) are aggregated using the absolute sum.
(a) Taxi Dataset
Group Name Mean Sum No° Features in Group Features (Importance)
Infra. (Built) 1048.92 9440.26 9 street type (3975.59), street surface (2310.95), number of street lanes (1738.58), max speed (580.54), longitude (341.46), distance city center (170.93), latitude (138.44), is street one way (93.78), parking (89.99)
Infra. (POI) 698.52 7683.71 11 count hospitals (1171.37), count hotels (1160.42), count shops (1017.01), count industry (834.12), count bars (752.56), count education (721.87), count restaurants (614.71), count subway lightrailway (485.64), count bus stops (443.02), count long-distance bus stops (259.23), count cafes (223.77)
Target Variable 288.51 288.51 1 target feature (288.51)
Connectivity 241.36 965.45 4 degree (318.97), closeness (240.02), betweenness (210.35), clustering (196.12)
Weather 165.26 2148.44 13 pressure (284.43), snowfall (271.74), hourly precipitation (246.82), temperature avg. (246.62), hourly wind speed (194.61), hourly relative humidity (193.12), temperature max (192.14), precipitation (140.94), temperature min (95.62), wind speed avg. (88.93), hourly dew point (88.49), hourly pressure (67.92), hourly temperature (37.05)
Features (Missing/Masked) 120.00 240.00 2 binary feature indicating missing (142.77), binary feature indicating masked (97.24)
Holidays 24.89 49.79 2 public holiday (32.41), school holiday (17.38)
Time 19.19 76.77 4 day of week (33.17), month (24.21), hour (13.93), weekend (5.46)
(b) Strava Dataset
Group Name Mean Sum No° Features in Group Features (Importance)
Time 1017.80 4071.21 4 month (1945.71), year (1143.46), day of week (789.43), weekend (192.60)
Infra. (POI) 481.10 2405.51 5 count education (771.11), count hotels (623.53), count hospitals (395.17), count shops (369.08), count industry (246.62)
Infra. (Built) 238.10 6428.72 27 street type (1920.03), street surface (1685.42), bicycle lane type (537.04), cyclability touring (467.14), max speed (330.28), is within cycling route (236.52), horticulture percent (210.71), cyclability commute (186.05), public facilities percent (168.03), arable land percent (115.82), latitude (103.64), number of street lanes (82.91), commercial area percent (54.69), grassland percent (51.94), weekend house area percent (51.58), water bodies percent (34.13), cyclability (27.73), longitude (26.18), distance city center (23.51), city square percent (22.88), residential use percent (22.62), waste disposal percent (17.19), allotment gardens percent (16.48), park area percent (16.45), forest area percent (10.17), traffic area percent (9.54), cemetery percent (0.04)
Features (Missing/Masked) 179.12 358.23 2 binary feature indicating missing (219.11), binary feature indicating masked (139.12)
Socioeconomic 165.45 2812.62 17 net migration per 100 (374.00), migration volume per 100 (361.02), gender distribution (307.00), share foreign eu nationals (252.13), greying index (232.83), old age dependency ratio (204.42), unemployment rate age 15 to 65 (157.57), share under 18 (152.26), share with migration background (118.03), share foreign non eu nationals (114.50), total fertility rate (111.94), youth dependency ratio (95.57), share foreign nationals (90.54), share 65+ (79.70), residents 5+ years same address (76.69), average age (50.31), total population (34.09)
Target Variable 86.47 86.47 1 target feature (86.47)
Connectivity 61.30 306.51 5 main cycling network (280.18), clustering (22.01), betweenness (1.86), closeness (1.50), degree (0.95)
Weather 40.20 361.79 9 temperature max (151.93), wind speed avg. (69.50), pressure (46.22), temperature avg. (35.51), snowfall (26.84), wind speed gust (13.31), temperature min (12.09), sunshine duration (3.34), precipitation (3.04)
Holidays 39.29 78.59 2 public holiday (72.79), school holiday (5.79)
Motorized 6.82 81.86 12 avg. speed cars (21.68), avg. speed all vehicles (20.40), avg. speed trucks (17.62), vehicle count all vehicles (4.85), avg. vehicle speed 6km (4.78), vehicle count trucks 6km (4.55), vehicle count cars (4.50), avg. speed cars 6km (1.21), vehicle count cars 6km (1.01), avg. speed trucks 6km (0.74), vehicle count all vehicles 6km (0.30), vehicle count trucks (0.22)
To improve interpretability and support future data collection, we conduct a feature importance analysis using integrated gradients (IG).
IG is not only compatible with graph neural networks but particularly well-suited to our setting, as its additive nature enables straightforward aggregation of attribution scores. Since GNNUI operates on temporal windows where the same features are repeated across multiple time steps per node, this property allows meaningful aggregation of attributions across both temporal (time window) and spatial (nodes) dimensions.
We compute IG using 50 interpolation steps and average the gradients along the path from baseline to input. Scores are then aggregated across time and space using the absolute sum. Finally, we merge one-hot encoded categorical features and features computed at multiple radii (e.g., hospital counts within 0.05–0.5 km).
Table 6 reports both individual and grouped feature importance scores (mean and sum) for the Strava and taxi data. We include group-level scores because many features are obtained together (e.g., weather variables), so identifying key feature groups helps guide future data collection. For both Strava and taxi data, infrastructure features dominate, including both built infrastructure and points of interest. Street type, street surface, and mode-specific infrastructure are particularly influential, with the number of lanes and maximum speed for taxis, and the bicycle lane type for Strava. However, the datasets show differences in their secondary drivers: for taxi data, the target variable and connectivity measures show medium importance, while for Strava, socioeconomic features and indicators for missing/masked data are relevant. Most notably, temporal features exhibit strikingly different importance patterns between the datasets. Time indicators are crucial for Strava (with month and year having the highest scores), but much less critical for taxi data. This difference likely reflects the distinct temporal characteristics of the datasets: taxi data consists of hourly observations across two months, while Strava data comprises daily observations across four years, capturing much broader temporal variation. Additionally, cycling behavior may inherently exhibit greater temporal variability than taxi usage, with seasonal patterns and recreational cycling trends creating more pronounced time-dependent effects.
It is important to note that these findings are closely tied to the estimation of Strava and taxi data specifically. While general tendencies are observable, the relative importance of features may differ when modeling overall cycling or motorized traffic volumes. This is especially relevant given the known biases in crowdsourced data. Strava users, for instance, tend to be younger, male, and potentially more influenced by amenities like restaurants than the average cyclist. For a broader feature importance analysis of general cycling patterns, refer to (Kaiser et al., 2025).
Appendix K Connectivity Measures
@@
Appendix L Combination of Adjacency Matrices
@@
We only explain the adapted architecture of GNNUi (with a usage of ZINB loss), as only this case is employed here. However, the procedure for the other losses is the same and can be transferred accordingly. For both adjacency matrices, we seperately proceed with the same procedure as if we were to use only one adjacency matrix, passing the input through the three DGCN layers, using a relu activation function after the first two and a linear one after the last one to obtain $H_{3}\in\mathbb{R}^{n_{o}+n_{m}\times z}$. We then concatenate those along the second dimension to $H_{3},c\in\mathbb{R}^{n_{o}+n_{m}\times 2z}$. This is then passed through a ZINB layer to obtain $n$, $p$, and $\pi$. This is depicted in Figure 7. Because the subnetworks for each adjacency matrix are trained separately up to H3, this approach results in a substantial increase in computational requirements.
@@
Table 7 presents the performance results of GNNUI, given the different combinations of adjacency matrices. Models that combine binary adjacency with another type perform well, likely benefiting from the strong performance of the binary structure. However, none of these combinations outperforms the standalone binary adjacency matrix, except for the KL for the taxi data (Binary is at 0.731 and Binary & Similarity at 0.744). Due to the small numerical difference, we think this is negligible. Interestingly, the combination of similarity and distance-based adjacency matrices improves upon the individual specifications in certain cases. For example, in New York, the Similarity + Distance Time combination achieves an MAE of 30.39, improving upon similarity (40.50) and distance time (31.68). While this suggests that hybrid approaches may be beneficial, they still fall short of the binary adjacency matrix. This may be because combining different adjacency matrices (e.g., binary and similarity) can introduce overlapping or weakly relevant connections, causing the model to be distracted by noisy or less informative edges and thereby diluting the strong local signal that the binary adjacency effectively captures.
@@
Appendix M Varying Percentage of Street Segments: For Similarity and Distance Based Adjacency Matrices